OpenStudy (anonymous):

How do I write this in summation notation? \[(\sin x) (\cos x)^{n - 1} + (\sin x) (\cos x)^{\frac{n - 1}{2}} + (\sin x) (\cos x)^{\frac{\frac{n - 1}{2}}{2}} + (\sin x) (\cos x)^{\frac{\frac{\frac{n - 1}{2}}{2}}{2}} ... \] n is an odd number and the series ends when (n-1)/2^k = 1, and the last term ends up being sin(x) only (no cos factor). For example, if n was 5, then 4 / 2^2 would be the last term in the series, so it looks like cos^1(x)sin(x) would be the last term, but the last term is actually sin(x) only. So if n was 5, the series would look like: \[(\sin x) (\cos x)^{4} + (\sin x) (\cos x)^{2} + (\sin x)\]

hartnn (hartnn):

i am just trying. the cos powers are 1,2,4,8,16.... so, 2k starts from 0 and ends to n-1 and factoring out sin x, \(\large \sum \limits_{k=0}^{n-1} \sin x \cos x^{2k}\)

hartnn (hartnn):

\(\large \sum \limits_{k=0}^{n-1} \sin x (\cos x)^{2k}\)

OpenStudy (anonymous):

If n = 5, then for your term you have cos^8(x)

hartnn (hartnn):

ohh...

hartnn (hartnn):

how about \(\large \sum \limits_{k=0}^{(n-1)/2} \sin x (\cos x)^{2k}\)

OpenStudy (anonymous):

That one would work, I guess. I'll post another problem

hartnn (hartnn):

okk..

OpenStudy (anonymous):

I got something slightly different. I used n = 2^m + 1 and set the sum over m

hartnn (hartnn):

i think we need to add restriction that k should be odd..

hartnn (hartnn):

and n as a power ? you'll get odd powers... m=1, n=3 m=2, n=5

OpenStudy (anonymous):

like if n = 5, then m = 2

hartnn (hartnn):

how about this : \(\large \sum \limits_{k=0}^{(n-1)/2} \sin x (\cos x)^{2^k}\)

OpenStudy (anonymous):

Oh

hartnn (hartnn):

2k will work, 2^k will not work..

OpenStudy (anonymous):

You are right